Circular No. 8869

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 R3 (GIBBS) A. R. Gibbs reports his discovery of another comet on unfiltered CCD

images obtained with the 1.5-m reflector in the course of the Mt. Lemmon Survey (“discovery” observation tabulated below); the four co-added 30-s frames show a well-condensed 8′′ coma and a faint, narrow tail 20′′ long in p.a. 240^{o}, while four co-added 80-s frames taken around Sept. 14.37 UT show a condensed 11′′ coma and a faint 30′′ tapered tail in p.a. 240^{o}. Four co-added 30-s exposures by Gibbs on Sept. 15.24–15.27 show a condensed 8′′ coma and a faint 10′′ tail in p.a. 225^{o}, while four co-added 80-s frames taken around Sept. 15.35–15.36 show a condensed 10′′ coma and a narrow 40′′ tail in p.a. 245^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometric observers have also commented on the object’s cometary appearance. G. Hug (Scranton, KS, 0.30-m reflector) writes that the object appeared diffuse on his images from Sept. 15.3. J. G. Ries (McDonald Observatory 0.76-m reflector) notes that her images from Sept. 15.4 show a coma and a faint tail ≈ 23′′ long toward the southwest. V. Gonano, E. Guido, and G. Sostero (Remanzacco, Italy, 0.45-m reflector) report that fourth co-added 60-s exposures taken around Sept. 16.0 show a diffuse coma nearly 15′′ in diameter with a broad tail nearly 25′′ long in p.a. 260^{o}. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 reflector) writes that his images from Sept. 16.0 show a 10′′ coma with a 30′′ tail in p.a. 240^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.14.29338  0^{h}41^{m}55.^{s}38  +1^{o}26′12′′.0  18.9

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-S02.

T = 2007 June 6.477 TT  ω = 292.^{o}699
e = 0.53777  Ω = 33.711 }2000.0
q = 2.20125 AU  i = 3.165
a = 4.76219 AU n^{o} = 0.094840 P = 10.4 years

COMET C/2007 N3 (LULIN) The following revised parabolic orbital elements for this comet (cf.

IAUC 8857) are from MPEC 2007-R56: T = 2009 Jan. 10.7452 TT, q = 1.211903 AU, ω = 136^{o}.8430, Ω = 338^{o}.4873, i = 178^{o}.3716 (equinox 2000.0).

2007 September 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green